OpenStudy (anonymous):

Prove sec x tan x (1-sin^2x)=___x

OpenStudy (lgbasallote):

what is that blank for? you have to prove and there's a blank?? o.O

OpenStudy (anonymous):

What is the proof

OpenStudy (anonymous):

\[\text{Sec}[x] \left(1-\text{Sin}[x]^2\right) \text{Tan}[x]\to \frac{1}{\text{Cos}[x]} \text{Cos}[x]^2 \frac{\text{Sin}[x]}{\text{Cos}[x]}\to \text{Sin}[x] \]